# Periodic Cosmic String Formation and Dynamics

### v1

@@ Appendix C Perturbative Stability Analysis at $\mathcal{J}_{0}\ll\mu v^{2}$ and $\bar{\zeta}_{0}\ll 1$
The term in the []delimited-[][\,$\cdots$][ ⋯ ] bracket must be set to zero in order to enforce the definition $\langle\zeta_{0}\rangle=0$, thereby implicitly defining $\bar{\rho}_{0}$ via $\bar{\rho}_{0}^{2}/v^{2}-\mathcal{J}_{0}^{2}/(\mu^{2}\bar{\rho}_{0}^{4})-1=0$; while this can actually be solved in closed form, the resulting expression is algebraically complicated so we omit it here. However, let us define
@@
$\zeta_{0}(\tilde{t})$ $=\bar{\zeta}_{0}\cos(\sqrt{2}\cdot\tilde{t}+\varphi_{\zeta})\>.$ (68)
@@
$\tilde{\omega}^{2}_{i}(\tilde{t})$ $=\tilde{k}^{2}-1+\bar{c}_{i}\left[1+\zeta_{0}(\tilde{t})\right]^{2}$ (70)
@@
$\partial_{\tilde{t}}^{2}\mathbb{X}_{i}+\tilde{\omega}_{i}^{2}\mathbb{X}_{i}$ $=0\>;$ $\tilde{\omega}_{i}^{2}$ $\equiv\frac{\omega_{i}^{2}}{\mu^{2}}\approx(\tilde{k}^{2}+c_{i}-1)+2c_{i}\zeta_{0}(\tilde{t})\>.$ (72)
As noted in the main text in the paragraph below (30), this casts the approximate EoM into the form of the Mathieu equation $\partial_{\tilde{t}}^{2}f+\tilde{\omega}_{i}^{2}(\tilde{t})\cdot f=0$ where $\tilde{\omega}_{i}^{2}(\tilde{t})\equiv\omega_{0,i}^{2}+2\epsilon_{i}\cos(\sqrt{2}(\tilde{t}-\tilde{t}_{0}))$, which exhibits both narrow and broad parametric resonance phenomena. For $\epsilon_{i}\ll\omega_{0,i}^{2}/2$, narrow resonance occurs Kovacic:2018weg : the primary narrow resonance exists for $1/2-\epsilon_{i}\lesssim\omega_{0,i}^{2}\lesssim 1/2+\epsilon_{i}$, while a secondary narrow resonance occurs when $62-\epsilon_{i}^{2}/6\lesssim\omega_{0,i}^{2}\lesssim 2+5\epsilon_{i}^{2}/$ 6; in fact, resonances exist for all $\omega_{0,i}^{2}\sim n^{2}/2$ for $n\in\mathbb{Z},n>0$ (the centers of these bands can also be offset by amounts $\propto\epsilon_{i}^{j}$ for some $j\in\mathbb{Z},j>0$), with widths $\Delta(\omega_{0}^{2})\propto\epsilon_{i}^{n}$. However, the higher-$n$ resonances are increasingly inefficient at driving growth: for instance, at small $\epsilon_{i}$, it can be shown that the largest Floquet exponents [^40] The growing solution amplitude increases by $\text{exp}[\tilde{\Gamma}\tilde{T}]$ in each period $\Delta\tilde{t}=\tilde{T}=\pi\sqrt{2}$ of the $\tilde{\omega}_{i}^{2}$ oscillation. $\tilde{\Gamma}$ for the primary band ($n=1$) are $\tilde{\Gamma}\sim\epsilon_{i}/\sqrt{2}$ for $\omega_{0,i}^{2}=1/2$ (i.e., at the band center), while those in the secondary band ($n=2$) are $\tilde{\Gamma}\sim\epsilon_{i}^{2}/(4\sqrt{2})$ for $\omega_{0,i}^{2}=2+\epsilon_{i}^{2}/3$ (i.e., at the band center); see, e.g., MagnusWinkler ; Kovacic:2018weg for brief introductions to Floquet analysis. For $\epsilon_{i}\gtrsim\omega_{0,i}^{2}/2$, the resonance is instead broad: the field $f$ exhibits durations where $\tilde{\omega}^{2}_{i}(\tilde{t})<0$, yielding explosive field growth Kofman:1997yn (except in certain narrow bands of parameter space).
@@ D.1 Limiting Case of $\rho\gg v$
In the limit $\mathcal{J}_{0}\ll\mu v^{2}$ and $\rho\gg v$, from (45) we have $\ddot{\rho}_{0}+DH\dot{\rho}_{0}+\rho_{0}^{3}\approx 0$. Rescaling $t\equiv x^{3(D+1)/(D+3)}$ and $\rho_{0}\equiv a^{-D/3}f$, we find $f^{\prime\prime}+(3(D+1)/(D+3))^{2}f^{3}-2D(D-3)/(D+3)^{2}f/x^{2}=0$, where ${}^{\prime}\equiv d/dx$. But for $0<D\leq 3$, this is clearly an undamped anharmonic oscillator equation equation with positive (and, for $D\neq 3$, time-dependent) frequency squared, so $f$ just oscillates with roughly fixed amplitude: $f\propto a^{0}\times g$, where $g$ is an oscillatory function (which we will drop below); this is strictly true for $D=3$, and true at late time for $D=2$.
These facts confirm that, in amplitude, $\rho_{0}\propto a^{-D/3}$. For large $\rho\gg v$, $V_{\rho}\propto a^{-4D/3}$ follows immediately. From $\mathcal{J}_{0}\sim\rho_{0}^{2}\dot{\theta}_{0}\propto a^{-D}$, we then also have $\dot{\theta}_{0}\propto a^{-D/3}$, from which it follows that $k_{\theta}\propto a^{-4D/3}$. Finally, in $D=3$ and in $D=2$ at late times, the EoM for $f$ takes the form $f^{\prime\prime}+cf^{3}\approx 0$ for some constant $c$. Multiplying through by $f^{\prime}$ and integrating once, we have $(f^{\prime})^{2}/2+cf^{4}/4\sim\text{const.}$. Since $f\propto a^{0}$, we must then have $f^{\prime}\propto a^{0}$. Moreover, it is fairly straightforward to show that $\dot{\rho}\propto x^{-4D/(D+3)}f^{\prime}$ at late times, from which it follows that $\dot{\rho}\propto a^{-2D/3}$ and therefore $k_{\rho}\propto a^{-4D/3}$.
@@ D.2 Limiting Case of $\rho\sim v$ (for $\mathcal{J}_{0}\ll\mu v^{2}$)
In the limit $\rho\simeq v$ with $\mathcal{J}_{0}\ll\mu v^{2}$, different scalings obtain. To see this, set $\mathcal{J}_{0}=0$ and $\rho=v+\delta$ in (45), expand in powers of $\delta$, and keep only the lowest-order terms. This gives $\ddot{\delta}+DH\dot{\delta}+2\mu^{2}\delta\approx 0$. Then take $\delta\equiv a^{-D/2}f$, which yields $\ddot{f}+\left[2\mu^{2}+D/(D+1)^{2}/t^{2}\right]f\approx 0$, which is again an oscillator equation with a time-dependent frequency, which goes to a simple harmonic oscillator equation at late times. As a result, we expect a roughly fixed-amplitude oscillation: $f\propto a^{0}\times h$, where $h$ is an oscillatory function (which we omit in what follows). Moreover, at late times, we have (by a similar argument as in App. D.1) that $\dot{f}^{2}/2+\mu^{2}f^{2}\approx\text{const.}$., which implies that we must have $f^{\prime}\propto a^{0}$ in amplitude too. But $\mathcal{J}_{0}\approx v^{2}\dot{\theta}\propto a^{-D}\Rightarrow\dot{\theta}\propto a^{-D}$. Putting this all together, it follows that $V\propto\mu^{2}\delta^{2}\propto a^{-D}$, $k_{\rho}\propto\dot{\rho}^{2}\propto a^{-D}\dot{f}^{2}\propto a^{-D}$ (at late times), and $k_{\theta}\propto v^{2}\dot{\theta}^{2}\propto a^{-2D}$.
@@ Appendix E String Burst Periodicity for General $\rho/v$
Note that, in the large $\rho/v$ limit, the term $-\mu^{2}\rho^{4}$ can also be dropped, which led to the $\rho\propto a^{-D/3}$ scaling we found in the main text at (47). However, as $\rho/v\rightarrow 1$, the $-\mu^{2}\rho^{4}$ term cannot be neglected, and the solution for $\rho(a)$ must be found from (82). [^42] Given the structure of (82), which is a cubic equation for $\rho^{2}$, this can in principle be done in closed form, but the expression is algebraically complicated, so we omit it here. Correspondingly, $\dot{\theta}(a)$ inside the counter-rotating regions can then be obtained with $\dot{\theta}(a)=j_{0}/[\rho(a)]^{2}$. The spacing $\Delta a$ between sequential string bursts (“inter-burst spacing”) can then be determined as a function of $a$ directly from (48), yielding

### v2

@@ Appendix C Perturbative Stability Analysis at $\mathcal{J}_{0}\ll\mu v^{2}$ and $\bar{\zeta}_{0}\ll 1$
The term in the $\cdots$] bracket must be set to zero in order to enforce the definition $\langle\zeta_{0}\rangle=0$, thereby implicitly defining $\bar{\rho}_{0}$ via $\bar{\rho}_{0}^{2}/v^{2}-\mathcal{J}_{0}^{2}/(\mu^{2}\bar{\rho}_{0}^{4})-1=0$; while this can actually be solved in closed form, the resulting expression is algebraically complicated so we omit it here. However, let us define
@@
[…]
@@
[…]
@@
[…]
As noted in the main text in the paragraph below (30), this casts the approximate EoM into the form of the Mathieu equation $\partial_{\tilde{t}}^{2}f+\tilde{\omega}_{i}^{2}(\tilde{t})\cdot f=0$ where $\tilde{\omega}_{i}^{2}(\tilde{t})\equiv\omega_{0,i}^{2}+2\epsilon_{i}\cos(\sqrt{2}(\tilde{t}-\tilde{t}_{0}))$, which exhibits both narrow and broad parametric resonance phenomena. For $\epsilon_{i}\ll\omega_{0,i}^{2}/2$, narrow resonance occurs Kovacic:2018weg : the primary narrow resonance exists for $1/2-\epsilon_{i}\lesssim\omega_{0,i}^{2}\lesssim 1/2+\epsilon_{i}$, while a secondary narrow resonance occurs when $2-\epsilon_{i}^{2}/6\lesssim\omega_{0,i}^{2}\lesssim 2+5\epsilon_{i}^{2}/6$; in fact, resonances exist for all $\omega_{0,i}^{2}\sim n^{2}/2$ for $n\in\mathbb{Z},n>0$ (the centers of these bands can also be offset by amounts $\propto\epsilon_{i}^{j}$ for some $j\in\mathbb{Z},j>0$), with widths $\Delta(\omega_{0}^{2})\propto\epsilon_{i}^{n}$. However, the higher-$n$ resonances are increasingly inefficient at driving growth: for instance, at small $\epsilon_{i}$, it can be shown that the largest Floquet exponents [^43] The growing solution amplitude increases by $\text{exp}[\tilde{\Gamma}\tilde{T}$] in each period $\Delta\tilde{t}=\tilde{T}=\pi\sqrt{2}$ of the $\tilde{\omega}_{i}^{2}$ oscillation. $\tilde{\Gamma}$ for the primary band ($n=1$) are $\tilde{\Gamma}\sim\epsilon_{i}/\sqrt{2}$ for $\omega_{0,i}^{2}=1/2$ (i.e., at the band center), while those in the secondary band ($n=2$) are $\tilde{\Gamma}\sim\epsilon_{i}^{2}/(4\sqrt{2})$ for $\omega_{0,i}^{2}=2+\epsilon_{i}^{2}/3$ (i.e., at the band center); see, e.g., MagnusWinkler ; Kovacic:2018weg for brief introductions to Floquet analysis. For $\epsilon_{i}\gtrsim\omega_{0,i}^{2}/2$, the resonance is instead broad: the field $f$ exhibits durations where $\tilde{\omega}^{2}_{i}(\tilde{t})<0$, yielding explosive field growth Kofman:1997yn (except in certain narrow bands of parameter space).
@@ D.1 Limiting Case of $\rho\gg v$
In the limit $\mathcal{J}_{0}\ll\mu v^{2}$ and $\rho\gg v$, from (45) we have $\ddot{\rho}_{0}+DH\dot{\rho}_{0}+\rho_{0}^{3}\approx 0$. Rescaling $t\equiv x^{3(D+1)/(D+3)}$ and $\rho_{0}\equiv a^{-D/3}f$, we find $f^{\prime\prime}+(3(D+1)/(D+3))^{2}f^{3}-2D(D-3)/(D+3)^{2}f/x^{2}=0$, where $^{\prime}\equiv d/dx$. But for $0<D\leq 3$, this is clearly an undamped anharmonic oscillator equation with positive (and, for $D\neq 3$, time-dependent) frequency squared, so $f$ just oscillates with roughly fixed amplitude: $f\propto a^{0}\times g$, where $g$ is an oscillatory function (which we will drop below); this is strictly true for $D=3$, and true at late time for $D=2$.
These facts confirm that, in amplitude, $\rho_{0}\propto a^{-D/3}$. For large $\rho\gg v$, $V_{\rho}\propto a^{-4D/3}$ follows immediately. From $\mathcal{J}_{0}\sim\rho_{0}^{2}\dot{\theta}_{0}\propto a^{-D}$, we then also have $\dot{\theta}_{0}\propto a^{-D/3}$, from which it follows that $k_{\theta}\propto a^{-4D/3}$. Finally, in $D=3$ and in $D=2$ at late times, the EoM for $f$ takes the form $f^{\prime\prime}+cf^{3}\approx 0$ for some constant $c$. Multiplying through by $f^{\prime}$ and integrating once, we have $(f^{\prime})^{2}/2+cf^{4}/4\sim\text{const.}$ Since $f\propto a^{0}$, we must then have $f^{\prime}\propto a^{0}$. Moreover, it is fairly straightforward to show that $\dot{\rho}\propto x^{-4D/(D+3)}f^{\prime}$ at late times, from which it follows that $\dot{\rho}\propto a^{-2D/3}$ and therefore $k_{\rho}\propto a^{-4D/3}$.
@@ D.2 Limiting Case of $\rho\sim v$ (for $\mathcal{J}_{0}\ll\mu v^{2}$)
In the limit $\rho\simeq v$ with $\mathcal{J}_{0}\ll\mu v^{2}$, different scalings obtain. To see this, set $\mathcal{J}_{0}=0$ and $\rho=v+\delta$ in (45), expand in powers of $\delta$, and keep only the lowest-order terms. This gives $\ddot{\delta}+DH\dot{\delta}+2\mu^{2}\delta\approx 0$. Then take $\delta\equiv a^{-D/2}f$, which yields $\ddot{f}+\left[2\mu^{2}+D/(D+1)^{2}/t^{2}\right]f\approx 0$, which is again an oscillator equation with a time-dependent frequency, which goes to a simple harmonic oscillator equation at late times. As a result, we expect a roughly fixed-amplitude oscillation: $f\propto a^{0}\times h$, where $h$ is an oscillatory function (which we omit in what follows). Moreover, at late times, we have (by a similar argument as in App. D.1) that $\dot{f}^{2}/2+\mu^{2}f^{2}\approx\text{const.}$, which implies that we must have $f^{\prime}\propto a^{0}$ in amplitude too. But $\mathcal{J}_{0}\approx v^{2}\dot{\theta}\propto a^{-D}\Rightarrow\dot{\theta}\propto a^{-D}$. Putting this all together, it follows that $V\propto\mu^{2}\delta^{2}\propto a^{-D}$, $k_{\rho}\propto\dot{\rho}^{2}\propto a^{-D}\dot{f}^{2}\propto a^{-D}$ (at late times), and $k_{\theta}\propto v^{2}\dot{\theta}^{2}\propto a^{-2D}$.
@@ Appendix E String Burst Periodicity for General $\rho/v$
Note that, in the large $\rho/v$ limit, the term $\mu^{2}\rho^{4}$ can also be dropped, which led to the $\rho\propto a^{-D/3}$ scaling we found in the main text at (47). However, as $\rho/v\rightarrow 1$, the $\mu^{2}\rho^{4}$ term cannot be neglected, and the solution for $\rho(a)$ must be found from (82). [^45] Given the structure of (82), which is a cubic equation for $\rho^{2}$, this can in principle be done in closed form, but the expression is algebraically complicated, so we omit it here. Correspondingly, $\dot{\theta}(a)$ inside the counter-rotating regions can then be obtained with $\dot{\theta}(a)=j_{0}/[\rho(a)]^{2}$. The spacing $\Delta a$ between sequential string bursts (“inter-burst spacing”) can then be determined as a function of $a$ directly from (48), yielding